OpenStudy (anonymous):

find the arc length of f(x)= ln(sec(x)) on the interval [0, (pi/4)] around the x-axis.

zepdrix (zepdrix):

Hey Sam :) What part we stuck on?

zepdrix (zepdrix):

We take a curve and break it into a bunch of small pieces of "length" \(\Large\rm ds\). To find arc length, we simply add up all of those pieces.\[\Large\rm S=\int\limits ds\]\[\Large\rm S=\int\limits \sqrt{1+\left(\frac{dy}{dx}\right)^2}~dx\]

zepdrix (zepdrix):

We can go over that ds = (stuff) if you need. Otherwise, looks like we just need a derivative to get things rolling.

zepdrix (zepdrix):

\[\Large\rm f(x)=\ln(\sec x)\]\[\Large\rm f'(x)=?\]

zepdrix (zepdrix):

Oh `around the x-axis`. So this is a surface area problem?

OpenStudy (anonymous):

\[(1\div \sec x)*secxtanx\]

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{1}{\sec x}~\sec x \tan x\]Ok good :o simplify

OpenStudy (anonymous):

tanx

OpenStudy (anonymous):

dont I have to use the formula